OpenStudy (anonymous):

What are the solutions of the equation: z2 - 12z + 36 = 0?

OpenStudy (anonymous):

what you mean is you want to find the value of z correct?

OpenStudy (anonymous):

umm yes i think so.

OpenStudy (anonymous):

Here are the answers. A. -6, 6 B. 6, 6 C. 6, -6 D. -6, -6

OpenStudy (anonymous):

ok then what do you think should be our first step?

OpenStudy (anonymous):

ok so you are looking for a point. based on the answer choices

OpenStudy (anonymous):

z^2+12z+36 =(z+6)(z+6)

OpenStudy (anonymous):

Umm Im not sure.if i am or not.But the answer i got was b, but i dont know if its correct or not.

OpenStudy (anonymous):

ok yes this is a quadratic equation you didnt put the z^2 in the first equation.

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (anonymous):

Haha lol how do you know?

OpenStudy (anonymous):

not sure.

OpenStudy (anonymous):

not sure how to factor or not sure if im right about what it is?

OpenStudy (anonymous):

Im confused.

OpenStudy (anonymous):

ok what are you confused about?

OpenStudy (anonymous):

Wait my equation is correct. im just not sure on the answer.

OpenStudy (anonymous):

Solution (z+6)(z+6) = z=-6 z= -6 This should be your answer D. -6,-6

OpenStudy (anonymous):

Ok thanks. Im closing.

OpenStudy (anonymous):

@jmartinez44 @Kenziehulse im sorry but d is not correct

OpenStudy (anonymous):

Ok then what do you think it is?

OpenStudy (anonymous):

It is not?

OpenStudy (anonymous):

z^2-12z+36=0 (x-6)(x-6) (6,6) the b variable is negative B is your answer

OpenStudy (anonymous):

My bad I didn't see the -12z I'm very sorry

OpenStudy (anonymous):

its ok jmartinez44

OpenStudy (anonymous):

so my answer choice b was correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks now im closing.

OpenStudy (anonymous):

I opened up another.